OpenStudy (anonymous):

What are the solutions of 4x^2 – 3x – 5 = 0?

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

Look up how to use the quadratic formula or use wolfram alpha like I did to find the solutions which are x = (3-sqrt(89))/8 and x = (3+sqrt(89))/8 If you have problems with it in form me of what your problems are and I can help you

OpenStudy (anonymous):

My choices are.. A .. x=3+-sqrt71/6 .... B.. X= -3+-sqrt71/8....C..x=3+-sqrt89/8 ... D.. -3+-sqrt89/6

OpenStudy (anonymous):

Want some advice Learn how to use the quadratic formula as it has applications in chemistry and math and you wont be able to escape it. I gave you the solutions to the problem already I'm not telling you what letter You are only cheating yourself by not learning it trust me